# Euclid: A blue galaxy population and a brightest cluster galaxy in the making in a MaDCoWS2 galaxy cluster candidate††thanks: This paper is published on behalf of the Euclid Consortium

### 07 Mar 2026 v2

@@ 4.3 Finding other assembling BCGs
The discovery of an assembling proto-BCG in one of the two Euclid Deep Fields covered by MaDCoWS2 suggests that these objects might be common – a hypothesis that is further supported by the presence of similar merging BCGs in the literature. Yet, it is difficult to assess the prevalence of these systems because high-resolution near-infrared (e.g. this study; see also Kuiper et al., 2011; Coogan et al., 2023) or (sub)millimeter (Miller et al., 2018, see also Rotermund et al. 2021) observations are necessary to reveal the complex structures of proto-BCGs.
The large areas of the sky that will be observed by the Euclid, Nancy Grace Roman, and Vera C. Rubin should allow more systematic detections of merging proto-BCGs. In particular, the Euclid Wide Survey will cover approximately deg2 at the same depth as the Q1 data release (Euclid Collaboration: Scaramella et al., 2022; Euclid Collaboration: Mellier et al., 2025; Euclid Collaboration: Aussel et al., 2025). Thus, assuming that our single detection in the overlap between Euclid Q1 Data Release and MaDCoWS2 (deg2) is representative of the density of proto-BCGs on the sky, we would expect to find about 400 proto-BCGs in the completed Euclid Wide Survey.
@@ Acknowledgements.
We thank C.C. Steidel for providing an archival sensitivity function for MOSFIRE. A.T., A.H.G., S.A.S., S.T., D.S., P.R.M.E and B.M. acknowledge the support of NASA ROSES Grant 12-EUCLID11-0004. The work of D.S. and P.R.M.E. was carried out at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with the National Aeronautics and Space Administration (80NM0018D0004). B.M. acknowledges funding support from the Jet Propulsion Lab (JPL) Euclid project. This work has made use of the Euclid Quick Release Q1 data from the Euclid mission of the European Space Agency (ESA), 2025, https://doi.org/10.57780/esa-2853f3b. The Euclid Consortium acknowledges the European Space Agency and a number of agencies and institutes that have supported the development of Euclid, in particular the Agenzia Spaziale Italiana, the Austrian Forschungsförderungsgesellschaft funded through BMIMI, the Belgian Science Policy, the Canadian Euclid Consortium, the Deutsches Zentrum für Luft- und Raumfahrt, the DTU Space and the Niels Bohr Institute in Denmark, the French Centre National d’Etudes Spatiales, the Fundação para a Ciência e a Tecnologia, the Hungarian Academy of Sciences, the Ministerio de Ciencia, Innovación y Universidades, the National Aeronautics and Space Administration, the National Astronomical Observatory of Japan, the Netherlandse Onderzoekschool Voor Astronomie, the Norwegian Space Agency, the Research Council of Finland, the Romanian Space Agency, the Swiss Space Office (SSO) at the State Secretariat for Education, Research, and Innovation (SERI), and the United Kingdom Space Agency. A complete and detailed list is available on the Euclid web site (www.euclid-ec.org/consortium/community/). This project used data obtained with the Dark Energy Camera (DECam), which was constructed by the Dark Energy Survey (DES) collaboration. Funding for the DES Projects has been provided by the U.S. Department of Energy, the U.S. National Science Foundation, the Ministry of Science and Education of Spain, the Science and Technology Facilities Council of the United Kingdom, the Higher Education Funding Council for England, the National Center for Supercomputing Applications at the University of Illinois at Urbana-Champaign, the Kavli Institute of Cosmological Physics at the University of Chicago, Center for Cosmology and Astro-Particle Physics at the Ohio State University, the Mitchell Institute for Fundamental Physics and Astronomy at Texas AM University, Financiadora de Estudos e Projetos, Fundao Carlos Chagas Filho de Amparo, Financiadora de Estudos e Projetos, Fundao Carlos Chagas Filho de Amparo Pesquisa do Estado do Rio de Janeiro, Conselho Nacional de Desenvolvimento Cientfico e Tecnolgico and the Ministrio da Cincia, Tecnologia e Inovao, the Deutsche Forschungsgemeinschaft and the Collaborating Institutions in the Dark Energy Survey. The Collaborating Institutions are Argonne National Laboratory, the University of California at Santa Cruz, the University of Cambridge, Centro de Investigaciones Enrgeticas, Medioambientales y TecnolgicasMadrid, the University of Chicago, University College London, the DES-Brazil Consortium, the University of Edinburgh, the Eidgenssische Technische Hochschule (ETH) Zrich, Fermi National Accelerator Laboratory, the University of Illinois at Urbana-Champaign, the Institut de Cincies de l’Espai (IEEC/CSIC), the Institut de Fsica d’Altes Energies, Lawrence Berkeley National Laboratory, the Ludwig-Maximilians Universitt Mnchen and the associated Excellence Cluster Universe, the University of Michigan, the National Optical Astronomy Observatory, the University of Nottingham, the Ohio State University, the University of Pennsylvania, the University of Portsmouth, SLAC National Accelerator Laboratory, Stanford University, the University of Sussex, and Texas AM University. This work is based in part on observations made with the Spitzer Space Telescope, which was operated by the Jet Propulsion Laboratory, California Institute of Technology under a contract with NASA. This research has made use of the NASA/IPAC Infrared Science Archive, which is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology. Some of the data presented herein were obtained at Keck Observatory, which is a private 501(c)3 non-profit organization operated as a scientific partnership among the California Institute of Technology, the University of California, and the National Aeronautics and Space Administration. The Observatory was made possible by the generous financial support of the W. M. Keck Foundation. This research has made use of the SVO Filter Profile Service ‘Carlos Rodrigo,’ funded by MCIN/AEI/10.13039/501100011033/ through grant PID2023-146210NB-I00.References
@@
- The Bimodal Galaxy Color Distribution: Dependence on Luminosity and Environment. ApJL 615, pp. L101–L104. External Links: ISSN 0004-637X, Link, Document Cited by: §1.
@@
- How filaments of galaxies are woven into the cosmic web. Natur 380, pp. 603–606. External Links: ISSN 0028-0836, Link, Document Cited by: §1.
@@
- A massive quiescent galaxy at redshift 4.658. Natur 619 (7971), pp. 716–719 (en). Note: Publisher: Nature Publishing Group External Links: ISSN 1476-4687, Link, Document Cited by: §3.3.3.
@@
- Euclid preparation. III. Galaxy cluster detection in the wide photometric survey, performance and algorithm selection. A&A 627, pp. A23. External Links: Document, 1906.04707 Cited by: §1.
@@
- Evidence of Runaway Gas Cooling in the Absence of Supermassive Black Hole Feedback at the Epoch of Cluster Formation. ApJL 898, pp. L50. External Links: ISSN 0004-637X, Link, Document Cited by: §1, §4.2.3.
@@
- A Massive, Cooling-Flow-Induced Starburst in the Core of a Highly Luminous Galaxy Cluster. Natur 488 (7411), pp. 349–352. Note: Comment: 11 pages, 3 figures, 1 table. Supplemental material contains 15 additional pages. Published in Nature External Links: ISSN 0028-0836, 1476-4687, Link, Document Cited by: §4.2.3.
@@
- An HST/WFC3-UVIS View of the Starburst in the Cool Core of the Phoenix Cluster. ApJL 765, pp. L37. External Links: ISSN 0004-637X, Link, Document Cited by: §4.2.3.
@@
- The heating of gas in a galaxy cluster by X-ray cavities and large-scale shock fronts. Natur 433 (7021), pp. 45–47 (en). External Links: ISSN 1476-4687, Link, Document Cited by: §1.
@@
- A massive core for a cluster of galaxies at a redshift of 4.3. Natur 556, pp. 469–472. External Links: ISSN 0028-0836, Link, Document Cited by: §1, §4.2.1, §4.2.2, §4.2.2, §4.2, §4.3.
@@
- Directly imaging the cooling flow in the Phoenix cluster. Natur 638, pp. 360–364. External Links: ISSN 0028-0836, Link, Document Cited by: §4.2.3.
@@
- SVO Filter Profile Service Version 1.0. Note: IVOA Working Draft 15 October 2012 External Links: Document Cited by: §2.2.1.
@@
- Simulations of the formation, evolution and clustering of galaxies and quasars. Natur 435, pp. 629–636. External Links: ISSN 0028-0836, Link, Document Cited by: §1.
@@
- COSMOS-Web galaxy groups: Evolution of red sequence and quiescent galaxy fraction. arXiv e-prints, pp. arXiv:2509.08040. External Links: Document Cited by: §1.
@@
- Reversal of Fortune: Confirmation of an Increasing Star Formation-Density Relation in a Cluster at z = 1.62. ApJL 719, pp. L126–L129. External Links: ISSN 0004-637X, Link, Document Cited by: §1.
@@
- Spectroscopic confirmation of a mature galaxy cluster at a redshift of 2. Natur 577, pp. 39–41. External Links: ISSN 0028-0836, Link, Document Cited by: §3.1.
@@
- Turbulent Heating in Galaxy Clusters Brightest in X-rays. Natur 515 (7525), pp. 85–87. Note: Comment: published in Nature External Links: ISSN 0028-0836, 1476-4687, Link, Document Cited by: §1.

### 13 Mar 2026 v3

@@ 4.3 Finding other assembling BCGs
The discovery of an assembling proto-BCG in one of the two Euclid Deep Fields covered by MaDCoWS2 suggests that these objects might be common – a hypothesis that is further supported by the presence of similar merging BCGs in the literature. Yet it is difficult to assess the prevalence of these systems because high-resolution near-infrared (e.g. this study; see also Kuiper et al. 2011; Coogan et al. 2023) or (sub)millimetre (Miller et al. 2018, see also Rotermund et al. 2021) observations are necessary to reveal the complex structures of proto-BCGs.
The large areas of the sky that will be observed by the Euclid, Nancy Grace Roman, and Vera C. Rubin should allow more systematic detections of merging proto-BCGs. In particular, the Euclid Wide Survey will cover approximately deg2 at the same depth as the Q1 data release (Euclid Collaboration: Scaramella et al. 2022; Euclid Collaboration: Mellier et al. 2025; Euclid Collaboration: Aussel et al. 2025). Thus, assuming that our single detection in the overlap between Euclid Q1 Data Release and MaDCoWS2 (deg2) is representative of the density of proto-BCGs on the sky, we would expect to find about 400 proto-BCGs in the completed Euclid Wide Survey.
@@ Acknowledgements.
We thank C.C. Steidel for providing an archival sensitivity function for MOSFIRE. A.T., A.H.G., S.A.S., S.T., D.S., P.R.M.E and B.M. acknowledge the support of NASA ROSES Grant 12-EUCLID11-0004. The work of D.S. and P.R.M.E. was carried out at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with the National Aeronautics and Space Administration (80NM0018D0004). B.M. acknowledges funding support from the Jet Propulsion Lab (JPL) Euclid project. This work has made use of the Euclid Quick Release Q1 data from the Euclid mission of the European Space Agency (ESA), 2025, https://doi.org/10.57780/esa-2853f3b. The Euclid Consortium acknowledges the European Space Agency and a number of agencies and institutes that have supported the development of Euclid, in particular the Agenzia Spaziale Italiana, the Austrian Forschungsförderungsgesellschaft funded through BMIMI, the Belgian Science Policy, the Canadian Euclid Consortium, the Deutsches Zentrum für Luft- und Raumfahrt, the DTU Space and the Niels Bohr Institute in Denmark, the French Centre National d’Etudes Spatiales, the Fundação para a Ciência e a Tecnologia, the Hungarian Academy of Sciences, the Ministerio de Ciencia, Innovación y Universidades, the National Aeronautics and Space Administration, the National Astronomical Observatory of Japan, the Netherlandse Onderzoekschool Voor Astronomie, the Norwegian Space Agency, the Research Council of Finland, the Romanian Space Agency, the Swiss Space Office (SSO) at the State Secretariat for Education, Research, and Innovation (SERI), and the United Kingdom Space Agency. A complete and detailed list is available on the Euclid web site (www.euclid-ec.org/consortium/community/). This project used data obtained with the Dark Energy Camera (DECam), which was constructed by the Dark Energy Survey (DES) collaboration. Funding for the DES Projects has been provided by the U.S. Department of Energy, the U.S. National Science Foundation, the Ministry of Science and Education of Spain, the Science and Technology Facilities Council of the United Kingdom, the Higher Education Funding Council for England, the National Center for Supercomputing Applications at the University of Illinois at Urbana-Champaign, the Kavli Institute of Cosmological Physics at the University of Chicago, Center for Cosmology and Astro-Particle Physics at the Ohio State University, the Mitchell Institute for Fundamental Physics and Astronomy at Texas AM University, Financiadora de Estudos e Projetos, Fundao Carlos Chagas Filho de Amparo, Financiadora de Estudos e Projetos, Fundao Carlos Chagas Filho de Amparo Pesquisa do Estado do Rio de Janeiro, Conselho Nacional de Desenvolvimento Cientfico e Tecnolgico and the Ministrio da Cincia, Tecnologia e Inovao, the Deutsche Forschungsgemeinschaft and the Collaborating Institutions in the Dark Energy Survey. The Collaborating Institutions are Argonne National Laboratory, the University of California at Santa Cruz, the University of Cambridge, Centro de Investigaciones Enrgeticas, Medioambientales y TecnolgicasMadrid, the University of Chicago, University College London, the DES-Brazil Consortium, the University of Edinburgh, the Eidgenssische Technische Hochschule (ETH) Zrich, Fermi National Accelerator Laboratory, the University of Illinois at Urbana-Champaign, the Institut de Cincies de l’Espai (IEEC/CSIC), the Institut de Fsica d’Altes Energies, Lawrence Berkeley National Laboratory, the Ludwig-Maximilians Universitt Mnchen and the associated Excellence Cluster Universe, the University of Michigan, the National Optical Astronomy Observatory, the University of Nottingham, the Ohio State University, the University of Pennsylvania, the University of Portsmouth, SLAC National Accelerator Laboratory, Stanford University, the University of Sussex, and Texas AM University. This work is based in part on observations made with the Spitzer Space Telescope, which was operated by the Jet Propulsion Laboratory, California Institute of Technology under a contract with NASA. This research has made use of the NASA/IPAC Infrared Science Archive, which is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology. Some of the data presented herein were obtained at Keck Observatory, which is a private 501(c)3 non-profit organization operated as a scientific partnership among the California Institute of Technology, the University of California, and the National Aeronautics and Space Administration. The Observatory was made possible by the generous financial support of the W. M. Keck Foundation. This research has made use of the SVO Filter Profile Service ‘Carlos Rodrigo’, funded by MCIN/AEI/10.13039/501100011033/ through grant PID2023-146210NB-I00.References
@@
- The Bimodal Galaxy Color Distribution: Dependence on Luminosity and Environment. ApJ 615, pp. L101–L104. External Links: ISSN 0004-637X, Link, Document Cited by: §1.
@@
- How filaments of galaxies are woven into the cosmic web. Nat 380, pp. 603–606. External Links: ISSN 0028-0836, Link, Document Cited by: §1.
@@
- A massive quiescent galaxy at redshift 4.658. Nat 619 (7971), pp. 716–719 (en). External Links: ISSN 1476-4687, Link, Document Cited by: §3.3.3.
@@
- Euclid preparation. III. Galaxy cluster detection in the wide photometric survey, performance and algorithm selection. A&A 627, pp. A23. External Links: Document, 1906.04707 Cited by: §1, §1.
@@
- Evidence of Runaway Gas Cooling in the Absence of Supermassive Black Hole Feedback at the Epoch of Cluster Formation. ApJ 898, pp. L50. External Links: ISSN 0004-637X, Link, Document Cited by: §1, §4.2.3.
@@
- A Massive, Cooling-Flow-Induced Starburst in the Core of a Highly Luminous Galaxy Cluster. Nat 488 (7411), pp. 349–352. Note: Comment: 11 pages, 3 figures, 1 table. Supplemental material contains 15 additional pages. Published in Nature External Links: ISSN 0028-0836, 1476-4687, Link, Document Cited by: §4.2.3.
@@
- An HST/WFC3-UVIS View of the Starburst in the Cool Core of the Phoenix Cluster. ApJ 765, pp. L37. External Links: ISSN 0004-637X, Link, Document Cited by: §4.2.3.
@@
- The heating of gas in a galaxy cluster by X-ray cavities and large-scale shock fronts. Nat 433 (7021), pp. 45–47 (en). External Links: ISSN 1476-4687, Link, Document Cited by: §1.
@@
- A massive core for a cluster of galaxies at a redshift of 4.3. Nat 556, pp. 469–472. External Links: ISSN 0028-0836, Link, Document Cited by: §1, §4.2.1, §4.2.2, §4.2.2, §4.2, §4.3.
@@
- Directly imaging the cooling flow in the Phoenix cluster. Nat 638, pp. 360–364. External Links: ISSN 0028-0836, Link, Document Cited by: §4.2.3.
@@
- SVO Filter Profile Service Version 1.0. Note: IVOA Working Draft 15 October 2012https://doi.org/10.5479/ADS/bib/2012ivoa.rept.1015R External Links: Document Cited by: §2.2.1.
@@
- Simulations of the formation, evolution and clustering of galaxies and quasars. Nat 435, pp. 629–636. External Links: ISSN 0028-0836, Link, Document Cited by: §1.
@@
- COSMOS-Web galaxy groups: Evolution of red sequence and quiescent galaxy fraction. A&A 707, pp. A87. External Links: ISSN 0004-6361, Link, Document Cited by: §1.
@@
- Reversal of Fortune: Confirmation of an Increasing Star Formation-Density Relation in a Cluster at z = 1.62. ApJ 719, pp. L126–L129. External Links: ISSN 0004-637X, Link, Document Cited by: §1.
@@
- Spectroscopic confirmation of a mature galaxy cluster at a redshift of 2. Nat 577, pp. 39–41. External Links: ISSN 0028-0836, Link, Document Cited by: §3.1.
@@
- Turbulent Heating in Galaxy Clusters Brightest in X-rays. Nat 515 (7525), pp. 85–87. Note: Comment: published in Nature External Links: ISSN 0028-0836, 1476-4687, Link, Document Cited by: §1.